# Imperceptible Protection Against Style Imitation from Diffusion Models

### v1

@@ Abstract
Recent progress in diffusion models has profoundly enhanced the fidelity of image generation. However, this has raised concerns about copyright infringements. While prior methods have introduced adversarial perturbations to prevent style imitation, most are accompanied by the degradation of artworks’ visual quality. Recognizing the importance of maintaining this, we develop a visually improved protection method that preserves its protection capability. To this end, we create a perceptual map to identify areas most sensitive to human eyes. We then adjust the protection intensity guided by an instance-aware refinement. We also integrate a perceptual constraints bank to further improve the imperceptibility. Results show that our method substantially elevates the quality of the protected image without compromising on protection efficacy.
1 Introduction
@@
The groundbreaking advancements in large-scale diffusion models have transformed media creation workflows [43, 2, 27, 40, 45]. These can be further enhanced in usability when integrated with external modules that accept multi-modal inputs [62, 28, 38, 18, 21]. Such innovations have also been pivotal in the realm of art creation [22, 50, 1]. Nevertheless, generative AI, while undoubtedly beneficial, brings concerns about its potential misuse. When someone exploits these to replicate artworks without permission, it introduces significant risks of copyright infringement. This style imitation becomes a serious threat to artists [49, 60].
@@
To counteract style imitation, previous studies have introduced adversarial perturbation [13, 32] to artwork, transforming it into an adversarial example that can resist few-shot generation or personalization methods [52, 30, 29, 61, 66, 46, 65]. Specifically, building on the Stable Diffusion (SD) model [43], they iteratively optimize the protected (or perturbed) image to fool the SD network, guided by the gradients from the image encoder or denoising UNet.
@@
While existing studies are effective in preventing style imitation, they do not prioritize the protected image’s quality. They leave discernible traces (or artifacts) on the protected images due to the inherent nature of adversarial perturbations. Moreover, we observed that compared to adversarial attacks on classifiers [31, 7], style protection requires more intense and globally dispersed perturbations. Consequently, despite the commendable protection performance, prior works run the risk of severely degrading the original artwork’s fidelity, making it less practical for real-world applications. While moderating the strength of protection could mitigate this, it introduces a trade-off, often compromising protection performance and making it challenging to achieve satisfactory results.
Figure 1: Impasto preserves the authenticity of artworks by impeding style imitation while reducing visual artifacts in the protected images. (Right) Impasto shows a significant enhancement in the balance between protection efficacy and imperceptibility.
@@
To alleviate this, we propose IMperceptible Protection Against STyle imitatiOn (Impasto; Fig. 2). We design this upon the principle of perception-aware protection, which focuses on perturbing regions less discernible to humans. Although many adversarial attack methods restrict perturbations to small areas to maximize imperceptibility [37, 6, 7], they are ineffective for style protection since personalization methods can exploit the references from non-perturbed textures. To circumvent this, we instead adopt a soft restriction strategy by relaxing the harsh condition of sparse constraints. This applies protection to the entire image but with modulated intensities. To implement this, it is crucial to identify which areas are perceptually more noticeable when perturbations are introduced. For this purpose, we analyze various perceptual maps that are suitable for the style protection task. Then, we propose a method that combines such perceptual maps and refines them in an image-specific manner. Such an instance-wise refinement sets itself apart from previous protection methods as it offers a way to strike the optimal balance between imperceptibility and protection performance.
To further enhance imperceptibility, we use a perceptual constraint bank. We delve into multiple feature spaces, examining the pixel and the latent spaces of both LPIPS [63] and CLIP [42]. Previous methods have also adopted perceptual constraints [61, 49]. However, their approach, typically limited to employing only one or two constraints, does not fully capitalize on the potential of perceptual models. In contrast, we employ a perceptual constraint bank to effectively steer towards more enhanced imperceptibility. Additionally, we integrate a soft restriction within these constraints. Diverging from previous methods that apply constraints uniformly via spatial averaging, we modulate the spatial influence of constraints in areas less perceptible to humans, thereby achieving closer alignment with the nuances of the human visual system. Intuitively, it may not be surprising that imperceptibility can be improved with a constraint bank. However, surprisingly, our work is the first investigation to apply constraints across multiple spaces, thereby enhancing fidelity without sacrificing robustness.
@@
To the best of our knowledge, Impasto is the pioneering approach that prioritizes the protected image’s quality in the style protection task. We show the effectiveness of Impasto through a broad range of experiments. It achieves robust protection against style imitation and maintains visual fidelity in protected images (Fig. 1). Impasto also significantly improves the trade-off balance between protection efficacy and image quality (Fig. 1, right). The flexibility of Impasto is demonstrated by its successful applications to existing protection frameworks [30, 46, 52, 29]. Impasto also maintains resilience and generalizes well against a range of countermeasures and personalization techniques, performing on par with the baseline methods. Our key contributions are highlighted as:
  • We propose Impasto, which applies human visual perception principles to achieve subtle and effective style protection in diffusion models.
  • Impasto incorporates perception-aware protection and a perceptual constraints bank to realize imperceptible but effective style protection.
  • We validate the efficacy of Impasto through various experiments, despite its straightforward modulo design. It can be effortlessly adopted in any existing protection method, confirming its ability to be deployed in real applications.
Figure 2: Model overview. Given an artwork, we construct a perceptual map $\mathcal{M}$ by integrating multiple JNDs, $\mathbf{M}$. Impasto optimizes imperceptible perturbation $\delta$ using $\mathcal{M}$ and concurrently updates the refinement parameters $\omega$, all while being steered by SD’s guidance. Consequently, the protected artwork impedes malicious users from fine-tuning or generating credible imitations. Blue dashed arrows denote the gradient flow.
2 Background
2.1 Diffusion Models
@@
Diffusion models have risen to prominence for their capacity to produce high-quality images [16, 9, 45, 39, 2]. In AI-assisted art production, Stable Diffusion (SD) [43] is widely used because of its exceptional quality and efficiency. In the SD model, an input image $\mathbf{x}$ is projected into a latent code via an image encoder $\mathcal{E}$ such that $\mathbf{z}=\mathcal{E}(\mathbf{x})$. A decoder $\mathcal{D}$ reverts the latent code to the image domain, represented as $\mathbf{x}^{\prime}=\mathcal{D}(\mathbf{z}^{\prime})$. The diffusion model derives a modified latent code $\mathbf{z}^{\prime}$ by incorporating external factors $y$, such as text prompt or other modalities [62, 21]. The training objective for SD at timestep $t$ is defined as:
@@
Here, a denoising UNet $\epsilon_{\theta}$ reconstructs the noised latent code $\mathbf{z}_{t}$, given $t$ and a conditioning vector $c(y)$. Leveraging the power of diffusion models, recent studies have investigated personalizing SD with a given few reference images. For example, textual inversion-based methods [12, 53, 1] exploit the embedding space of CLIP [42] while freezing the denoising UNet. On the other hand, model optimization-based methods [44, 24, 50] directly update the UNet.
2.2 Protection Against Style Imitation
Previous protection methods introduce adversarial perturbations $\delta$ to image $\mathbf{x}$, making protected image $\hat{\mathbf{x}}=\mathbf{x}+\delta$ through projected gradient descent (PGD) [32], a renowned algorithm in the adversarial attack task.
Encoder-based methods [49, 46, 61, 29] update $\delta$ under the guidance of the VAE encoder $\mathcal{E}$. They aim to maximize the distance between the encoded feature of the original image $\mathbf{x}$ and protected image $\hat{\mathbf{x}}$. In practice, many methods instead minimize the distance between the protected image $\hat{\mathbf{x}}$ and target image $\mathbf{y}$:
While $L_{\infty}$ norm ($||\delta||_{\infty}\leq\eta$;  $\eta$ is a protection budget) is widely used constraints, GLAZE [49] adopts LPIPS [63] and DUAW [61] employs SSIM [55].
UNet-based methods [52, 30, 46, 65, 29] update $\delta$ under the guidance of the denoising UNet, $\epsilon_{\theta}$, maximizing diffusion loss, $\mathcal{L_{SD}}$ as:
Upon this, Anti-DreamBooth [52] integrates DreamBooth training and Mist [29] merges Eq. 2 and 3, enhancing performance and robustness in various scenarios.
Since Impasto is versatile and can be effortlessly integrated into any existing protection methods, we generalize the style protection as below formulation.
where $\mathcal{L_{SP}}=-\lambda_{\mathcal{E}}\mathcal{L_{E}}(\mathbf{x}+\delta,\mathbf{y})+\lambda_{\mathcal{SD}}\mathcal{L_{SD}}(\mathcal{E}(\mathbf{x}+\delta))$. Then, we employ PGD [32] to get a protected image $\hat{\mathbf{x}}$. Let $\mathbf{x}^{(0)}$ denote the original artwork. The protected image of $i$-th optimization step is generated by a signed gradient ascent with step function sgn and step length $\alpha$ as given by:
where $\Pi_{\mathcal{N}_{\eta}(\mathbf{x})}$ is the projection onto the $L_{\infty}$ neighborhood around $\mathbf{x}$ with radius $\eta$. This process is repeated $N$ steps as $\hat{\textbf{x}}=\mathbf{x}^{(N)}$.
2.3 Imperceptible Adversarial Examples
The concept of imperceptibility is an actively investigated subject in adversarial attacks; some studies target specific elements; e.g. the low-frequency components [14, 31]. Another approach use advanced constraints [48]; color distance [64] or quality assessment [54] are adopted. Several methods focus on restricting perturbation regions; leveraging $L_{0}$ norm to produce sparse perturbation [37, 6] or limiting perturbations to tiny salient regions [7]. Our research draws inspiration from these studies. However, attacking discriminative models is fundamentally distinct from that of targeting generative models. Hence, we employ a specialized strategy designed specifically for disrupting style imitation.
Protected GeneratedProtected GeneratedProtected Generated
Refer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to caption
(a) Original(b) Full protection(c) Partial protection
Figure 3: Sparse restriction in style protection. We compare the full image protection with partial one that simulates sparse restriction. For partial protection, we only apply perturbation in the facial region (no perturbation in red inlet). Generated result shows that partial protection is inadequate for style protection, as personalization methods can capitalize on unprotected areas when learning the artwork’s style.
3 Method— Impasto
3.1 Perception-Aware Protection (PAP)
Naive approach— Sparse restriction. In adversarial attacks, perturbations are often confined to a sparse region to increase imperceptibility [37, 6, 7]. Yet, we observed that such restriction does not sufficiently prevent style imitation (Fig. 3). In this analysis, we compare the protection encompassing the entire image (Fig. 3b) against the one that applies perturbation to facial region only (Fig. 3c). It is shown that the partial protection cannot protect original artwork against DreamBooth [44] and we argue that this is because the personalization method can leverage textures from unprotected regions. Overall, sparse restriction may confine the perturbation region too aggressively to be applied in style protection. Therefore, we relax such assumptions to better align with this task.
Soft restriction. To address the limitation of sparse restriction, we instead employ a soft restriction strategy. It protects the entire image but with varying intensities across different regions. To this end, we introduce a perceptual map, $\mathcal{M}\in\mathbb{R}^{d}$, where $d$ is the number of pixels of an image. This map reflects the human sensitivity to subtle alterations; a value near 1.0 indicates a region with the highest perceptibility, while a value close to 0.0 signifies a region where changes are hardest to notice. With this map $\mathcal{M}$, we define perception-aware protection loss $\mathcal{L_{PAP}}(\mathbf{x},\delta,\mathbf{y},\mathcal{M})$, which use soft restriction-based $L_{p}$ norm constraint as:
where * denotes element-wise multiplication. Employing a soft restriction-based $L_{p}$ norm controls perturbations to be suppressed in regions with high perceptual visibility to humans and amplified in areas with lower perceptibility. The subsequent objective is now to quantify perceptual sensitivity to distortions.
(a) Input (b) LA (c) CM (d) CSF (e) Std. (f) Entropy (g) Avg. (h) IWR
Figure 4: Examples of perceptual maps. We employ multiple JND models in our analysis. Darker region corresponds to increased protection intensity. The perceptual maps include luminance adaptation (LA), contrast masking (CM), contrast sensitivity function (CSF), standard deviation (Std.), and entropy. Impasto constructs perceptual map $\mathcal{M}$ by spatially averaging these estimations (g) or an learnable manner (h).
Perceptual map analysis. To build a perceptual map $\mathcal{M}$ in a simple yet effective manner, we investigate the just noticeable difference (JND) concept [58] inspired by the human visual system. JND represents the minimum intensity of stimulus (perturbation in our context) required to produce a noticeable change in visual perception. The intent behind the JND estimation model is to determine this perceptual threshold for every image pixel. Given that the fundamental premise of JND— to quantify human sensitivity to subtle changes— aligns with our objective of perception-aware protection, we focus on analyzing 1) the effectiveness of JNDs in our formulation, and 2) which JND models yield the best results. To this end, as depicted in Fig. 4, we compare the following JND models. Detailed explanations for each model are described in Suppl.
  • Luminance adaptation (LA): Perturbations are less visible in regions of very low or high luminance and more noticeable in moderate lighting conditions [19]. Hence, we modulate protection strength based on pixel luminance with a fixed adaptation model (Fig. 4b).
  • Contrast masking (CM): Perturbations can seamlessly blend into regions with intricate textures, while they leave distinct traces on flat surfaces. To simulate this, we utilize the luminance contrast (or change) of a region to measure the complexity of the pixel [26, 57] (Fig. 4c).
  • Contrast sensitivity function (CSF): Given the band-pass characteristics of the human visual system, we utilize a frequency-based JND model [56]. The human eye is receptive to signals at modulated frequencies while exhibiting insensitivity to high-frequency components. Consequently, perturbations overlaid on high-frequency signals (e.g. edges) are less perceptible (Fig. 4d).
  • Standard deviation: To assess the spatial structure of an image, we calculate the standard deviation of local image blocks, inspired by SSIM [55]. This measures the image’s structural complexity, which correlates with the sensitivity to subtle perturbations (Fig. 4e).
  • Entropy: The entropy of an image block is computed to quantify the amount of information or complexity within a local region [59] (Fig. 4f).
Table 1 presents the protection performance when adopting the JND-based perceptual map. Image quality is assessed through DISTS [10], and protection performance is evaluated using FID [15]. The baseline is trained with Eq. 5 and other models are trained via Eq. 6 with corresponding JNDs. For all the JNDs, we inverse them i.e. $M^{k}=1-\text{JND}^{k}$, since a higher JND threshold represents lower sensitivity to changes. When compared to the baseline, perceptual map improves image quality (DISTS) across the board, but it also leads to a compromise in protection performance (FID). Among the JNDs, LA demonstrates the best protection performance. We conjecture that in many artworks, the majority of areas fall high or low-luminance, thereby maintaining perturbations strength high across extensive regions. However, as illustrated in Fig. 4b, even simple textures like the sky can have strong perturbations, placing LA at the lower image quality. On the other hand, CSF, Std, and Entropy generally apply high perturbations only to specific areas, such as edges, resulting in most regions being not fully protected and consequently, causing a huge degradation in protection performance. CM’s protection intensity is also determined by spatial changes but this covers more detailed local regions (see fields in Fig. 4c) and also being based on contrast, leading to both high quality and effective protection.
Table 1: Perceptual map analysis. DISTS measures image quality, while FID evaluates protection performance.
Map DISTS ($\downarrow$) FID ($\uparrow$)
@@ 3.1 Perception-Aware Protection (PAP)
Perceptual map. In our earlier investigation, we observed that the JND model fits surprisingly well with soft restriction approach, offering an effective way for the imperceptible style protection. Consequently, Impasto employs a perceptual map constructed using JND estimates, considering its simplicity and effectiveness. Among JNDs we analyzed (Table 1), LA and CM emerge as superior in the quality-protection trade-off. However, these results represent the average scores across a dataset and we note that some images exhibit better trade-off results with different JND models. Especially, since artworks have diverse styles, the best combinations can differ; for example, in Fig. 4, Std. and entropy demonstrate best performance for top image while bottom image shows a preference for CM and CSF. In addition, in practical scenarios, users seeking to protect their artwork are soley concerned with their specific pieces. Hence, it is crucial to ensure that the protection method is effectively applied to every individual artwork.
@@
Therefore, we apply all the JNDs listed in Table 1 simultaneously to create the perceptual map, since relying on a single JND could result in decreased performance for some artworks. Formally, for an artwork $\mathbf{x}$ requiring protection, we initially generate a corresponding perceptual map $\mathcal{M}$ with a collection of JNDs, $\mathbf{M}=\{M^{i},\dots,M^{K}\}$, where $K$ is the number of JNDs. To integrate multiple JNDs, the simplest method involves using a spatial average: $\mathcal{M}=\frac{1}{K}\textstyle\sum_{k=1}^{K}M^{k}$. As we will demonstrate in the experimental results, this straightforward approach proves to be a surprisingly effective universal algorithm for creating perceptual map. Considering its simplicity and efficacy, we adopt spatially averaged JNDs to construct the initial perceptual map.
@@
Instance-wise refinement. Although the above method is more effective in many scenarios than the single JND, applying a uniform averaged map across all artworks may still be suboptimal for some artworks. Moreover, the optimal perceptual map corresponding to an artwork varies depending on both artwork’s structure and the applied protective perturbations. Even for identical artwork, the detectability of perturbations is affected by changing these as human sensitivity varies with distortion type [11]. To this end, we propose an instance-wise refinement (IWR), customizing the perceptual map $\mathcal{M}$ for each artwork. During optimization, the perceptual map is refined through a weighted sum: $\mathcal{M}(\omega)=\textstyle\sum_{k=1}^{K}\text{softmax}(\mathbf{\omega})^{k}*M^{k}$, where $\mathbf{\omega}$ is a set of learnable parameters that adjust the contributions of each JND to $\mathcal{M}$. The refinement parameters $\mathbf{\omega}$ is optimized using the objective function below:
$\mathcal{L_{M}}=||\mathcal{L_{SP}}(\mathbf{x}+\mathcal{M}^{\prime}\odot\delta)-\mathcal{L_{SP}}(\mathbf{x}+\mathcal{M}(\omega)\odot\delta)||^{2}_{2}+\left(\textstyle\sum_{i=1}^{d}|\mathcal{M}(\omega)_{i}*\delta_{i}|^{p}\right)^{\frac{1}{p}}$ (7)
@@
with $\mathcal{M}^{\prime}$ being the initial perceptual map before the refinement steps. In Eq. 7, the former term enforces the consistency of the refined perceptual map $\mathcal{M}$ with the initial map $\mathcal{M}^{\prime}$ by minimizing the discrepancy in protection loss between them. The latter term enhances the perception-aware protection for a given specific image. Algorithm 1 (L7-10) details the procedure for this refinement. As demonstrated in Fig. 4h, instance-wise refinement has a pronounced effect as compared to the naive averaging approach (Fig. 4g), better capturing the nuances of image-specific perceptual sensitivity.
Data: Image $\mathbf{x}$, target image $\mathbf{y}$, all perceptual maps $\mathbf{M}$, step length $\alpha$

### v2

@@ Abstract
Recent progress in diffusion models has profoundly enhanced the fidelity of image generation, but it has raised concerns about copyright infringements. While prior methods have introduced adversarial perturbations to prevent style imitation, most are accompanied by the degradation of artworks’ visual quality. Recognizing the importance of maintaining this, we introduce a visually improved protection method while preserving its protection capability. To this end, we devise a perceptual map to highlight areas sensitive to human eyes, guided by instance-aware refinement, which refines the protection intensity accordingly. We also introduce a difficulty-aware protection by predicting how difficult the artwork is to protect and dynamically adjusting the intensity based on this. Lastly, we integrate a perceptual constraints bank to further improve the imperceptibility. Results show that our method substantially elevates the quality of the protected image without compromising on protection efficacy.
Introduction
@@
The groundbreaking advancements in large-scale diffusion models have transformed media creation workflows (Rombach et al. 2022; Balaji et al. 2022; Li et al. 2022; Podell et al. 2023; Saharia et al. 2022). These can be further enhanced in usability when integrated with external modules that accept multi-modal inputs (Zhang, Rao, and Agrawala 2023; Li et al. 2023; Mou et al. 2023; Huang et al. 2023; Kim et al. 2023). Such innovations have also been pivotal in the realm of art creation (Ko et al. 2023; Sohn et al. 2023; Ahn et al. 2023). Nevertheless, generative AI, while undoubtedly beneficial, brings concerns about its potential misuse. When someone exploits these to replicate artworks without permission, it introduces significant risks of copyright infringement. This style imitation becomes a serious threat to artists (Shan et al. 2023; Xiang 2022).
@@
To counteract style imitation, previous studies have introduced adversarial perturbation (Goodfellow, Shlens, and Szegedy 2014; Madry et al. 2017) to artwork, transforming it into an adversarial example that can resist few-shot generation or personalization methods (Van Le et al. 2023; Liang et al. 2023; Liang and Wu 2023; Ye et al. 2023; Zheng et al. 2023; Salman et al. 2023; Zhao et al. 2023). Specifically, building on the Stable Diffusion (SD) model (Rombach et al. 2022), they iteratively optimize the protected (or perturbed) image to fool the diffusion models, guided by the gradients from the image encoder or denoising UNet.
@@
While existing studies are effective in preventing style imitation, they do not prioritize the protected image’s quality. They leave discernible traces (or artifacts) on the protected images due to the inherent nature of adversarial perturbations. Moreover, we observed that unlike adversarial attacks on classifiers (Luo et al. 2022; Dai et al. 2023), style protection requires more intense and globally dispersed perturbations. Consequently, despite the commendable protection performance, prior works run the risk of severely degrading the original artwork’s fidelity, making it less practical for real-world applications. While moderating the strength of protection could mitigate this, it introduces a trade-off, often compromising protection performance and making it challenging to achieve satisfactory results.
@@
To address this, we propose Impasto (Figure 1). We design upon the principle of perception-aware protection, which focuses on perturbing regions less discernible to humans. While many studies in the adversarial attacks regime limit perturbations to small areas to maximize imperceptibility (Croce and Hein 2019; Dai et al. 2023), they are ineffective for style protection, as personalization methods can leverage references from non-perturbed textures. We instead adopt a soft restriction strategy, which protect the entire image but with modulated intensities, relaxing the harsh condition of sparse constraints. To implement this, it is crucial to identify which areas are perceptually more noticeable when perturbations are introduced. For this purpose, we analyze various perceptual maps that are suitable for our task. Then, we propose a method that combines such perceptual maps and refines them in an image-specific manner. Such an instance-wise refinement sets itself apart from previous protection methods as it offers a way to strike the optimal balance between imperceptibility and protection performance.
For better imperceptibility, we propose a difficulty-aware protection that automatically adjusts the protection strength based on how easily the area can be protected. To achieve this, we compute a difficulty map, which predicts the protection difficulty of a given artwork. The difficulty map is then combined with the perceptual map to determine the final protection intensity. Lastly, we introduce a perceptual constraint bank, which uses various features to apply perceptual constraints. It may seem unsurprising that a constraint bank can improve imperceptibility intuitively. However, our study is the first to apply constraints across multiple spaces, improving fidelity without compromising robustness.
Figure 1: Model overview. (a) Previous methods protect image through an iterative process that updates perturbations $\delta$ with PGD. (b) On this basis, Impasto constructs a perceptual map $\mathcal{M}$ and a difficulty map $\mathcal{M_{D}}$ to adaptively refine the perturbation for imperceptibility. Additionally, Impasto employs a constraint bank in the PGD process to achieve better imperceptibility. Red-bordered boxes indicate trainable parameters. IWR denotes instance-wise refinement.
@@
To the best of our knowledge, Impasto is the pioneering approach that prioritizes the protected image’s quality in the style protection task. It achieves robust protection against style imitation and maintains visual fidelity in protected images thus it can significantly improve the trade-off balance between protection efficacy and image quality. Moreover, Impasto is versatile, enabling its attachment to any existing protection frameworks (Liang et al. 2023; Salman et al. 2023; Van Le et al. 2023; Liang and Wu 2023). Impasto also maintains resilience and generalizes well against a range of countermeasures and personalization techniques, performing on par with the baseline methods.
Background
@@
Diffusion Models. These have gain prominence for their capacity to produce high-quality images (Ho, Jain, and Abbeel 2020; Dhariwal and Nichol 2021; Saharia et al. 2022; Nichol et al. 2021; Balaji et al. 2022). In AI-assisted art production, Stable Diffusion (SD) (Rombach et al. 2022) is widely used because of its efficiency. In SD, an input image $\mathbf{x}$ is projected into a latent code via an image encoder $\mathcal{E}$; $\mathbf{z}=\mathcal{E}(\mathbf{x})$. A decoder $\mathcal{D}$ reverts the latent code to the image domain; $\mathbf{x}^{\prime}=\mathcal{D}(\mathbf{z}^{\prime})$. The diffusion model derives a modified code $\mathbf{z}^{\prime}$ using external factors $y$, such as text prompt or other modalities (Zhang, Rao, and Agrawala 2023; Kim et al. 2023). The training objective for SD at timestep $t$ is defined as:
@@
Here, a denoising UNet $\epsilon_{\theta}$ reconstructs the noised latent code $\mathbf{z}_{t}$, given $t$ and a conditioning vector $c(y)$.
(a) No protect (b) Partial protect (c) Full protect
Figure 2: Comparison of restriction types. We compare partial and full protection by fine-tuning diffusion models using protected (or original) images via DreamBooth (Ruiz et al. 2023). For partial protection, perturbations are applied to the center (or facial) region. The results indicate that partial protection is insufficient for style protection, as personalization methods can utilize unprotected areas to learn the artwork’s style. Reference images are shown in Suppl.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Input LA CM CSF Stdev Entropy Average IWR
Config. DISTS ($\downarrow$) FID ($\uparrow$)
@@
Figure 3: Examples of perceptual maps. In this study, we use luminance adaptation (LA), contrast masking (CM), contrast sensitivity function (CSF), standard deviation (Stdev), and entropy. (Left) Perceptual Map visualization. Darker region corresponds to increased protection intensity. Impasto constructs perceptual map $\mathcal{M}$ by spatially averaging these estimations (Avgerage) or an learnable manner (IWR). (Right) Perceptual map comparison. We evaluate the protection performance of each perceptual map. Image quality is assessed through DISTS (Ding et al. 2020), and protection performance is evaluated via FID (Heusel et al. 2017). The baseline is trained with Eq. 5 and other models are trained via Eq. 6 with corresponding JNDs.
Protection Against Style Imitation. Previous methods introduce adversarial perturbations $\delta$ to image $\mathbf{x}$, making protected image $\hat{\mathbf{x}}=\mathbf{x}+\delta$ through projected gradient descent (PGD) (Madry et al. 2017). 1) Encoder-based methods (Shan et al. 2023; Salman et al. 2023; Ye et al. 2023; Liang and Wu 2023) (e.g. PhotoGuard) update $\delta$ under the guidance of the encoder $\mathcal{E}$. They aim to maximize the distance between the feature of the original and protected image. In practice, many methods minimize the distance between the protected image and target image $\mathbf{y}$ instead:
While $L_{\infty}$ norm ($||\delta||_{\infty}\leq\eta$;  $\eta$ is a budget) is widely used, GLAZE (Shan et al. 2023) adopts LPIPS (Zhang et al. 2018) and DUAW (Ye et al. 2023) employs SSIM (Wang et al. 2004). 2) Diffusion-based methods (Van Le et al. 2023; Liang et al. 2023; Salman et al. 2023; Zhao et al. 2023; Liang and Wu 2023) (e.g. AdvDM) update $\delta$ under the guidance of UNet, $\epsilon_{\theta}$, maximizing diffusion loss $\mathcal{L_{SD}}$:
Upon this, Anti-DreamBooth (Van Le et al. 2023) integrates DreamBooth training and Mist (Liang and Wu 2023) merges Eq. 2 and 3, enhancing performance and robustness in various scenarios. Since Impasto is versatile and can be effortlessly integrated into any existing protection methods, we generalize the style protection as below formulation.
where $\mathcal{L_{SP}}=-\lambda_{\mathcal{E}}\mathcal{L_{E}}(\mathbf{x}+\delta,\mathbf{y})+\lambda_{\mathcal{SD}}\mathcal{L_{SD}}(\mathcal{E}(\mathbf{x}+\delta))$. Then, we employ PGD to get a protected image $\hat{\mathbf{x}}$. Let $\mathbf{x}^{(0)}$ denote the original image. The protected image of $i$-th optimization step is generated by a signed gradient ascent with step function sgn and step length $\alpha$ as given by:
where $\Pi_{\mathcal{N}_{\eta}(\mathbf{x})}$ is the projection onto $L_{\infty}$ neighborhood with radius $\eta$. This process is repeated $N$ steps as $\hat{\textbf{x}}=\mathbf{x}^{(N)}$.
Method
Perception-Aware Protection
Naive sparse restriction. In adversarial attacks, perturbations are often confined to a sparse region to increase imperceptibility (Croce and Hein 2019; Dai et al. 2023). Yet, we observed that they can not sufficiently prevent style imitation. In Figure 2, we compare the protection encompassing the entire image against the one that applies perturbation to facial region only. It is shown that the partial protection cannot protect original artwork against DreamBooth (Ruiz et al. 2023) and we argue that this is because the personalization method can leverage textures from unprotected regions. Overall, sparse restriction may confine the perturbed region too aggressively, thus we relax such assumptions to better align with the style protection task.
Soft restriction. In this study, we instead adopt a soft restriction strategy. It protects the entire image but with varying intensities across different regions. To this end, we introduce a perceptual map, $\mathcal{M}\in\mathbb{R}^{d}$, where $d$ is the number of pixels of an image. This map reflects the human sensitivity to subtle alterations; a value near 1.0 indicates a region with the highest perceptibility, while a value close to 0.0 signifies a region where changes are hardest to notice. With $\mathcal{M}$, we define perception-aware protection loss $\mathcal{L_{PAP}}(\mathbf{x},\delta,\mathbf{y},\mathcal{M})$, which use soft restriction-based $L_{p}$ norm constraint as:
where * denotes element-wise multiplication. Employing a soft restriction-based $L_{p}$ norm controls perturbations to be suppressed in regions with high perceptual visibility to humans and amplified in areas with lower visibility. The question now is, how can we determine perceptual sensitivity?
Perceptual map. To build a perceptual map $\mathcal{M}$ in a simple yet effective manner, we investigate the just noticeable difference (JND) concept (Wu, Shi, and Lin 2019) inspired by the human visual system. JND represents the minimum intensity of stimulus (perturbation in our context) required to produce a noticeable change in visual perception. The intent behind the JND estimation model is to determine this perceptual threshold. Given that the fundamental premise of JND— to quantify human sensitivity to subtle changes— aligns with our objective of perception-aware protection, we focus on analyzing the effectiveness of JNDs in our formulation, and which JND models yield the good results.
To this end, as shown in Figure 3, we compare five widely used JNDs: 1) luminance adaptation (LM), which considers pixel luminance (Jarsky et al. 2011), 2) contrast masking (CM), which measures scene complexity by luminance contrast (Legge and Foley 1980; Wu et al. 2013a), 3) contrast sensitivity function (CSF), which utilizes frequency signals (Wei and Ngan 2009), 4) standard deviation (Stdev), which measures spatial structure with patch-wise statistics, and 5) entropy, which computes complexity with patch-wise entropy. More details are described in Suppl.
@@
In our observation, most JND models fit surprisingly well with soft restriction (Figure 3, right). Consequently, we employ a JND-based perceptual maps, considering its simplicity and effectiveness. Note that among JNDs we used, LA and CM emerge as superior in the quality-protection trade-off. However, these results are the average scores across a dataset and we note that some artworks exhibit better trade-off with other JNDs. Especially, since artworks have diverse styles, the best combinations can differ; for example, in Figure 3 (left), Stdev and entropy are best for top image while bottom image shows a preference for CM and CSF. Moreover, in practical scenarios, users who protect their artwork are solely concerned with their specific pieces, not the average scores. Thus, it is crucial to ensure that the protection method is effectively applied to every individual artwork.
@@
We therefore apply all the JNDs simultaneously to create the perceptual map, since relying on a single JND could result in decreased performance for some artworks. Formally, for an artwork $\mathbf{x}$ requiring protection, we initially generate a corresponding perceptual map $\mathcal{M}$ with a collection of JNDs, $\mathbf{M}=\{M^{i},\dots,M^{K}\}$, where $K$ is the number of JNDs. To integrate multiple JNDs, the simplest method involves using a spatial average: $\mathcal{M}=\frac{1}{K}\textstyle\sum_{k=1}^{K}M^{k}$.
@@
Instance-wise refinement. Although the above method is more effective in many scenarios than the single JND, applying a uniform averaged map across all artworks may still be suboptimal for some artworks. Moreover, the optimal perceptual map corresponding to an artwork varies depending on both artwork’s structure and the applied protective perturbations. Even for identical artwork, the detectability of perturbations is affected by changing these as human sensitivity varies with distortion type (Dodge and Karam 2017). Considering these, we propose an instance-wise refinement (IWR), customizing the perceptual map $\mathcal{M}$ for each artwork. During optimization steps, the perceptual map is refined through a weighted sum: $\mathcal{M}(\omega)=\textstyle\sum_{k=1}^{K}\text{softmax}(\mathbf{\omega})^{k}*M^{k}$, where $\mathbf{\omega}$ is a set of learnable parameters that adjust the contributions of each JND to $\mathcal{M}$. The refinement parameters $\mathbf{\omega}$ is optimized using the objective function below:
$\mathcal{L_{M}}$ $=||\mathcal{L_{SP}}(\mathbf{x}+\mathcal{M}^{\prime}\odot\delta)-\mathcal{L_{SP}}(\mathbf{x}+\mathcal{M}(\omega)\odot\delta)||^{2}_{2}$
$+\left(\textstyle\sum_{i=1}^{d}|\mathcal{M}(\omega)_{i}*\delta_{i}|^{p}\right)^{\frac{1}{p}},$ (7)
@@
with $\mathcal{M}^{\prime}$ being the initial perceptual map before the refinement steps. In Eq. 7, the former term enforces the consistency of the refined perceptual map $\mathcal{M}$ with the initial map $\mathcal{M}^{\prime}$ by minimizing the discrepancy in protection loss between them. The latter term enhances the perception-aware protection for a given specific image. As shown in Figure 3, instance-wise refinement (IWR) has a pronounced effect as compared to the naive averaging, better capturing the nuances of image-specific perceptual sensitivity.
Figure 4: Protection difficulty. We present the protection efficacy of PhotoGuard (Salman et al. 2023) across 20 artworks (x-axis). Notably, the level of style protection varies widely between artworks, thus the average score (red dashed line) is less meaningful to users who protect their artwork.
Data: Image $\mathbf{x}$, target image $\mathbf{y}$, perceptual maps $\mathbf{M}$